# Scaling the Authoring of AutoTutors with Large Language Models

### 14 Feb 2024 v1

@@ Abstract.
Large Language Models (LLMs) have found several use cases in education, ranging from automatic question generation to essay evaluation. In this paper, we explore the potential of using Large Language Models (LLMs) to author Intelligent Tutoring Systems. A common pitfall of LLMs is their straying from desired pedagogical strategies such as leaking the answer to the student, and in general, providing no guarantees. We posit that while LLMs with certain guardrails can take the place of subject experts, the overall pedagogical design still needs to be handcrafted for the best learning results. Based on this principle, we create a sample end-to-end tutoring system named MWPTutor, which uses LLMs to fill in the state space of a pre-defined finite state transducer. This approach retains the structure and the pedagogy of traditional tutoring systems that has been developed over the years by learning scientists but brings in additional flexibility of LLM-based approaches. Through a human evaluation study on two datasets based on math word problems, we show that our hybrid approach achieves a better overall tutoring score than an instructed, but otherwise free-form, GPT-4. MWPTutor is completely modular and opens up the scope for the community to improve its performance by improving individual modules or using different teaching strategies that it can follow.
@@ 2.2. Large Language Models
Large Language Models (LLMs) are Transformer-based models that serve as general-purpose tools in natural language processing (Bommasani et al., 2022). With recent improvements in LLMs and the availability of LLM-as-a-service models like GPT (OpenAI, 2023) and BARD (Yarkin and Coon, 2022), it has become easy to use LLMs as tutors by prompting them with the desired qualities of a tutor and then letting them interact with a student. Some works in the programming domain like GPTutor (Chen et al., 2023) and TestGPT (Phung et al., 2023) attempt to use GPT4 similar to information delivery systems. There have also been systems designed to function as Dialog agents, like Ruffle&Riley (Schmucker et al., 2023) which includes multiple GPT4 agents functioning with different prompts, and SPOCK (Sonkar et al., 2023) which attempts to distill out tutoring knowledge from GPT4 into a smaller LLM. A notable commercial usage of a LLM Tutor is seen in Khanmingo,222khanacademy.org/khan-labs which uses the GPT4 backend for an interactive tutor.
@@ 4.4. Answer and Path Detection
To execute the pedagogical strategies described above, we need to be able to detect if the student has given a correct answer. Further, the hint state should only trigger if the student fails to propose a correct follow-up. For this, we need to have a way to match paths. We run into the same matching problems described in Section 4.2, and resort to similar basic techniques. The correct answer detection algorithm is shown in Algorithm 2. For states other than pump, it takes a single reference step and checks if its RHS is present in the student utterance. For the pump state, all possible followups of the last step are passed as reference steps. If it can find the RHS of any of these in the student utterance, it marks that as the next step and proceeds to the next step. If not, it considers a path match to have occurred if every operand on the left-hand side of some step is present. If multiple steps fulfill this, or no step does, we pick the one with the best BertScore alignment to the student utterance.
@@ 5.1. Setup
To implement MWPTutor, we require a dataset with math word problems and their solutions, as well as incorrect student solutions. We make use of the test sets of two datasets, namely MathDial (Macina et al., 2023) and MultiArith. MultiArith consists of simple Math Word Problems that are usually solvable in 1 to 3 steps. GPT3 was already able to solve over 90% of these problems with the right prompts (Kojima et al., 2023), and in our tests, we found that GPT4 was able to get a perfect score in the MWP solver setting. With this dataset, we want to test if we can achieve the same performance in the interactive tutoring setting. This dataset does not provide any wrong solutions needed to instantiate the student model, so we make use of errors made by a MetaMath solver as the original student solution where possible, and leave the remaining questions as “No idea how to begin”.
@@ 8. Conclusion
These issues can be fixed by better LLMs and more sophisticated techniques outlined in future work. Several potential steps have already been outlined in Section 4, and we hope to soon have better versions of MWPTutor and other similar systems that can realize one-on-one tutoring at scale.
@@ Acknowledgements.
- Bommasani et al. (2022) Rishi Bommasani, Drew A. Hudson, Ehsan Adeli, Russ Altman, Simran Arora, Sydney von Arx, Michael S. Bernstein, Jeannette Bohg, Antoine Bosselut, Emma Brunskill, Erik Brynjolfsson, Shyamal Buch, Dallas Card, Rodrigo Castellon, Niladri Chatterji, Annie Chen, Kathleen Creel, Jared Quincy Davis, Dora Demszky, Chris Donahue, Moussa Doumbouya, Esin Durmus, Stefano Ermon, John Etchemendy, Kawin Ethayarajh, Li Fei-Fei, Chelsea Finn, Trevor Gale, Lauren Gillespie, Karan Goel, Noah Goodman, Shelby Grossman, Neel Guha, Tatsunori Hashimoto, Peter Henderson, John Hewitt, Daniel E. Ho, Jenny Hong, Kyle Hsu, Jing Huang, Thomas Icard, Saahil Jain, Dan Jurafsky, Pratyusha Kalluri, Siddharth Karamcheti, Geoff Keeling, Fereshte Khani, Omar Khattab, Pang Wei Koh, Mark Krass, Ranjay Krishna, Rohith Kuditipudi, Ananya Kumar, Faisal Ladhak, Mina Lee, Tony Lee, Jure Leskovec, Isabelle Levent, Xiang Lisa Li, Xuechen Li, Tengyu Ma, Ali Malik, Christopher D. Manning, Suvir Mirchandani, Eric Mitchell, Zanele Munyikwa, Suraj Nair, Avanika Narayan, Deepak Narayanan, Ben Newman, Allen Nie, Juan Carlos Niebles, Hamed Nilforoshan, Julian Nyarko, Giray Ogut, Laurel Orr, Isabel Papadimitriou, Joon Sung Park, Chris Piech, Eva Portelance, Christopher Potts, Aditi Raghunathan, Rob Reich, Hongyu Ren, Frieda Rong, Yusuf Roohani, Camilo Ruiz, Jack Ryan, Christopher Ré, Dorsa Sadigh, Shiori Sagawa, Keshav Santhanam, Andy Shih, Krishnan Srinivasan, Alex Tamkin, Rohan Taori, Armin W. Thomas, Florian Tramèr, Rose E. Wang, William Wang, Bohan Wu, Jiajun Wu, Yuhuai Wu, Sang Michael Xie, Michihiro Yasunaga, Jiaxuan You, Matei Zaharia, Michael Zhang, Tianyi Zhang, Xikun Zhang, Yuhui Zhang, Lucia Zheng, Kaitlyn Zhou, and Percy Liang. 2022. On the Opportunities and Risks of Foundation Models. arXiv:2108.07258 [cs.LG]
@@
- Kojima et al. (2023) Takeshi Kojima, Shixiang Shane Gu, Machel Reid, Yutaka Matsuo, and Yusuke Iwasawa. 2023. Large Language Models are Zero-Shot Reasoners. arXiv:2205.11916 [cs.CL]
@@
- Macina et al. (2023) Jakub Macina, Nico Daheim, Sankalan Pal Chowdhury, Tanmay Sinha, Manu Kapur, Iryna Gurevych, and Mrinmaya Sachan. 2023. MathDial: A Dialogue Tutoring Dataset with Rich Pedagogical Properties Grounded in Math Reasoning Problems. arXiv:2305.14536 [cs.CL]
@@
- Phung et al. (2023) Tung Phung, Victor-Alexandru Pădurean, José Cambronero, Sumit Gulwani, Tobias Kohn, Rupak Majumdar, Adish Singla, and Gustavo Soares. 2023. Generative AI for Programming Education: Benchmarking ChatGPT, GPT-4, and Human Tutors. In Proceedings of the 2023 ACM Conference on International Computing Education Research - Volume 2 (Chicago, IL, USA) (ICER ’23). Association for Computing Machinery, New York, NY, USA, 41–42. https://doi.org/10.1145/3568812.3603476

### 27 Feb 2024 v2

@@ Abstract.
Large Language Models (LLMs) have found several use cases in education, ranging from automatic question generation to essay evaluation. In this paper, we explore the potential of using LLMs to author Intelligent Tutoring Systems. A common pitfall of using LLMs as tutors is their straying from desired pedagogical strategies such as leaking the answer to the student, and in general, providing no guarantees on the validity or appropriateness of the tutor assistance. We argue that while LLMs with certain guardrails can take the place of subject experts, the overall pedagogical design still needs to be handcrafted for the best learning results. Based on this principle, we create a sample end-to-end tutoring system named MWPTutor, which uses LLMs to fill in the state space of a pre-defined finite state transducer. This approach retains the structure and the pedagogy of traditional tutoring systems that has been developed over the years by learning scientists but brings in additional flexibility of LLM-based approaches. Through a human evaluation study on two datasets with math word problems, we show that our hybrid approach achieves a better overall tutoring score than an instructed, but otherwise free-form, GPT-4. MWPTutor is completely modular and opens up the scope for the community to improve its performance by refining its individual modules or using different teaching strategies that it can follow.
@@ 2.2. Large Language Models
Large Language Models (LLMs) are Transformer-based models that serve as general-purpose tools in natural language processing (Bommasani et al., 2021). With recent improvements in LLMs and the availability of LLM-as-a-service models like GPT (OpenAI, 2023) and BARD (Yarkin and Coon, 2022), it has become easy to use LLMs as tutors by prompting them with the desired qualities of a tutor and then letting them interact with a student. Some works in the programming domain like GPTutor (Chen et al., 2023) and Phung et al. (Phung et al., 2023b, a) attempt to use GPT4 similar to information delivery systems. There have also been systems designed to function as Dialog agents, like Ruffle&Riley (Schmucker et al., 2023) which includes multiple GPT4 agents functioning with different prompts, and SPOCK (Sonkar et al., 2023) which attempts to distill out tutoring knowledge from GPT4 into a smaller LLM. A notable commercial usage of a LLM Tutor is seen in Khanmingo,222khanacademy.org/khan-labs which uses the GPT4 backend for an interactive tutor.
@@ 4.4. Answer and Path Detection
To execute the pedagogical strategies described above, we need to detect if the student has given a correct answer. Further, the hint state should only trigger if the student fails to propose a correct follow-up. For this, we need to have a way to match paths. We run into the same matching problems described in Section 4.2, and resort to similar basic techniques. The correct answer detection algorithm is shown in Algorithm 2. For states other than pump, it takes a single reference step and checks if its RHS is present in the student utterance. For the pump state, all possible followups of the last step are passed as reference steps. If it can find the RHS of any of these in the student utterance, it marks that as the next step and proceeds to the next step. If not, it considers a path match to have occurred if every operand on the left-hand side of some step is present. If multiple steps fulfill this, or no step does, we pick the one with the best BertScore alignment to the student utterance.
@@ 5.1. Setup
To implement MWPTutor, we require a dataset with math word problems and their solutions, as well as incorrect student solutions. We make use of the test sets of two datasets, namely MathDial (Macina et al., 2023) and MultiArith. MultiArith consists of simple Math Word Problems that are usually solvable in 1 to 3 steps. GPT3 was already able to solve over 90% of these problems with the right prompts (Kojima et al., 2022), and in our tests, we found that GPT4 was able to get a perfect score in the MWP solver setting. With this dataset, we want to test if we can achieve the same performance in the interactive tutoring setting. This dataset does not provide any wrong solutions needed to instantiate the student model, so we make use of errors made by a MetaMath solver as the original student solution where possible, and leave the remaining questions as “No idea how to begin”.
@@ 8. Conclusion
These issues can be fixed by better LLMs and more sophisticated techniques outlined in future work. Several potential steps have already been outlined, and we hope to soon have better versions of MWPTutor and other similar systems that can realize one-on-one tutoring at scale.
@@ Acknowledgements.
- Bommasani et al. (2021) Rishi Bommasani, Drew A. Hudson, Ehsan Adeli, Russ Altman, Simran Arora, Sydney von Arx, Michael S. Bernstein, Jeannette Bohg, Antoine Bosselut, Emma Brunskill, Erik Brynjolfsson, S. Buch, Dallas Card, Rodrigo Castellon, Niladri S. Chatterji, Annie S. Chen, Kathleen A. Creel, Jared Davis, Dora Demszky, Chris Donahue, Moussa Doumbouya, Esin Durmus, Stefano Ermon, John Etchemendy, Kawin Ethayarajh, Li Fei-Fei, Chelsea Finn, Trevor Gale, Lauren E. Gillespie, Karan Goel, Noah D. Goodman, Shelby Grossman, Neel Guha, Tatsunori Hashimoto, Peter Henderson, John Hewitt, Daniel E. Ho, Jenny Hong, Kyle Hsu, Jing Huang, Thomas F. Icard, Saahil Jain, Dan Jurafsky, Pratyusha Kalluri, Siddharth Karamcheti, Geoff Keeling, Fereshte Khani, O. Khattab, Pang Wei Koh, Mark S. Krass, Ranjay Krishna, Rohith Kuditipudi, Ananya Kumar, Faisal Ladhak, Mina Lee, Tony Lee, Jure Leskovec, Isabelle Levent, Xiang Lisa Li, Xuechen Li, Tengyu Ma, Ali Malik, Christopher D. Manning, Suvir P. Mirchandani, Eric Mitchell, Zanele Munyikwa, Suraj Nair, Avanika Narayan, Deepak Narayanan, Benjamin Newman, Allen Nie, Juan Carlos Niebles, Hamed Nilforoshan, J. F. Nyarko, Giray Ogut, Laurel Orr, Isabel Papadimitriou, Joon Sung Park, Chris Piech, Eva Portelance, Christopher Potts, Aditi Raghunathan, Robert Reich, Hongyu Ren, Frieda Rong, Yusuf H. Roohani, Camilo Ruiz, Jack Ryan, Christopher R’e, Dorsa Sadigh, Shiori Sagawa, Keshav Santhanam, Andy Shih, Krishna Parasuram Srinivasan, Alex Tamkin, Rohan Taori, Armin W. Thomas, Florian Tramèr, Rose E. Wang, William Wang, Bohan Wu, Jiajun Wu, Yuhuai Wu, Sang Michael Xie, Michihiro Yasunaga, Jiaxuan You, Matei A. Zaharia, Michael Zhang, Tianyi Zhang, Xikun Zhang, Yuhui Zhang, Lucia Zheng, Kaitlyn Zhou, and Percy Liang. 2021. On the Opportunities and Risks of Foundation Models. ArXiv (2021). https://crfm.stanford.edu/assets/report.pdf
@@
- Kojima et al. (2022) Takeshi Kojima, Shixiang (Shane) Gu, Machel Reid, Yutaka Matsuo, and Yusuke Iwasawa. 2022. Large Language Models are Zero-Shot Reasoners. In Advances in Neural Information Processing Systems, S. Koyejo, S. Mohamed, A. Agarwal, D. Belgrave, K. Cho, and A. Oh (Eds.), Vol. 35. Curran Associates, Inc., 22199–22213. https://proceedings.neurips.cc/paper_files/paper/2022/file/8bb0d291acd4acf06ef112099c16f326-Paper-Conference.pdf
@@
- Macina et al. (2023) Jakub Macina, Nico Daheim, Sankalan Pal Chowdhury, Tanmay Sinha, Manu Kapur, Iryna Gurevych, and Mrinmaya Sachan. 2023. MathDial: A Dialogue Tutoring Dataset with Rich Pedagogical Properties Grounded in Math Reasoning Problems. In The 2023 Conference on Empirical Methods in Natural Language Processing. https://openreview.net/forum?id=fyza2OQ9NI
@@
- Phung et al. (2023a) Tung Phung, José Cambronero, Sumit Gulwani, Tobias Kohn, Rupak Majumdar, Adish Singla, and Gustavo Soares. 2023a. Generating High-Precision Feedback for Programming Syntax Errors using Large Language Models. In Proceedings of the 16th International Conference on Educational Data Mining, Mingyu Feng, Tanja Käser, and Partha Talukdar (Eds.). International Educational Data Mining Society, Bengaluru, India, 370–377. https://doi.org/10.5281/zenodo.8115653
- Phung et al. (2023b) Tung Phung, Victor-Alexandru Pădurean, José Cambronero, Sumit Gulwani, Tobias Kohn, Rupak Majumdar, Adish Singla, and Gustavo Soares. 2023b. Generative AI for Programming Education: Benchmarking ChatGPT, GPT-4, and Human Tutors. In Proceedings of the 2023 ACM Conference on International Computing Education Research - Volume 2 (Chicago, IL, USA) (ICER ’23). Association for Computing Machinery, New York, NY, USA, 41–42. https://doi.org/10.1145/3568812.3603476